find tan3A-tan2A-tanA

find


tan3A-tan2A-tanA

Grade:12

1 Answers

askiitianexpert arulmani
14 years ago
tan 3A - tan 2A - tan A tan 3A = tan (A + 2A) = (tan A + tan 2A) / (1 - tan A tan 2A) tan 3A = (tan A + tan 2A) / (1 - tan A tan 2A) Bring denominator of Right hand side (2) to the left => tan 3A * (1 - tan A tan 2A) = (tan A + tan 2A) => tan 3A - tan A tan 2A tan 3A = tan A + tan 2A => tan 3A - tan 2A - tan A = tan A tan 2A tan 3A Solution is : tan 3A - tan 2A - tan A = tan A tan 2A tan 3A